I'm so lost and this is due tomorrow, if anyone can help that would be great!
Find all solutions of the equation?
3 cos(theta) + 3 = 2 sin^2(theta)
Thanks again!

Answer by Alan3354(69443)


3 cos(t) + 3 = 2 sin^2(t)
3 cos(t) + 3 = 2*(1 - cos^2(t))
2cos^2(t) + 3cos(t) + 1 = 0
(2cos + 1)*(cos + 1) = 0
cos = -1 --> theta = pi + 2n*pi, n = 0,1,2,3...
---------
cos = -1/2 --> theta = 2pi/3, 4pi/3 + 2n*pi, n = 0,1,2,3...